# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

### v1

@@ Abstract
Accurate material characterization and model calibration are essential for computationally-supported high-consequence engineering decisions. Current characterization and calibration methods (1) use simplified test specimen geometries and global data, (2) cannot guarantee that sufficient characterization data is collected for a specific model of interest, (3) use deterministic methods that provide best-fit parameter values with no uncertainty quantification, and (4) are sequential, inflexible, and time-consuming.
This work brings together several recent advancements into an improved workflow called Interlaced Characterization and Calibration (ICC) that advances the state-of-the-art in constitutive model calibration. The ICC paradigm (1) employs tools to efficiently use full-field data to calibrate high-fidelity material models, (2) aligns the data needed with the data collected by adopting an optimal experimental design protocol, (3) quantifies parameter uncertainty through Bayesian inference, and (4) incorporates these advances into a quasi real-time feedback loop. The ICC framework is demonstrated here on the calibration of a material model using simulated full-field data for an aluminum cruciform specimen being deformed bi-axially. The cruciform is actively driven through the myopically optimal load path using Bayesian optimal experimental design, which selects load steps that yield the maximum expected information gain (EIG). To aid in numerical stability and preserve computational resources, the full-field data is dimensionally reduced via principal component analysis (PCA), and fast surrogate models which approximate the input-output relationships of the expensive finite element model are used. The tools developed and demonstrated here show that high-fidelity constitutive models can be efficiently and reliably calibrated with quantified uncertainty, thus supporting credible decision-making and potentially increasing the agility of solid mechanics modeling by enabling utilization of computational simulations at earlier stages of the design cycle.SAND2024-15320O
@@
Bayesian optimal experimental design , load path optimization , uncertainty quantification , constitutive model calibration , digital image correlation , principal component analysis , Gaussian Process surrogate modeling
@@ 1 Introduction
Computational simulation is relied upon to understand complex engineering problems and provides essential information in decision-making. In finite element analysis of solid mechanics, constitutive relationships describe the response of a material to various external stimuli. Typically, these constitutive equations contain material-dependent parameters that are determined through material characterization and model calibration. Calibration and subsequent validation of the constitutive equations and parameters is an essential part of accurate computational simulations and structural analyses. Thus, there is a corresponding level of emphasis placed on the calibration process, both in terms of efficiency and accuracy, in order to establish confidence in predictions of the engineering problem at hand.
@@
Traditional characterization and calibration procedures are often suboptimal in terms of efficiency and accuracy, which challenges the effective use of computational simulation for design purposes. The two processes (material characterization and model calibration) are typically performed sequentially and independent of each other, and it is not uncommon for multiple experimental campaigns to be needed in order to obtain the proper data for calibration [1]. In addition to the procedure, the type of data collected is an important consideration for calibration efficiency. Traditionally, tests which target specific stress states (i.e., uniaxial tension, notched tension, top hat shear, etc.) have been used for calibration. These tests provide limited information per specimen; therefore, large test matrices are needed which include various tests, sample geometries and cut directions in order to reveal complex material behaviors. As a result of the current characterization and calibration workflow in tandem with traditional testing procedures, the model calibration process is elongated and expensive, thus delaying the support of computational simulation for design and engineering decisions.
@@
Several recent advancements have led to improvements in constitutive model calibration. For one, the use of digital image correlation (DIC) to collect rich full-field data through camera-based measurements has led to a paradigm shift in material characterization. Also known by the community as Material Testing 2.0 (MT2.0) [2], full-field DIC data collected from complex testing configurations (specimen shape plus loading) has greatly improved the information content that can be garnered from a single experiment for model calibration [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Moreover, researchers have developed several different methods—including Chebychev and Zernicke polynomials [14, 15, 16, 17, 18, 19] and various types of wavelets [20, 21, 22]—for decomposing the field data into a spectral domain in order to reduce the dimensionality of the data while retaining its rich information content. However, selection of the proper experimental configuration, or design, which greatly influences the suitability of the data for model calibration, remains a challenging task. Thus, the second area of improvement for model calibration is the development of procedures for optimal experimental design within the context of solid mechanics. In general, there are two different approaches which have been adopted: 1) optimization of the geometry of the specimen [23, 24, 25, 10], or 2) optimization of the load path of deformation [26, 27]. Both approaches use various techniques to construct an experiment which will yield the most useful data for model calibration, and, consequently, improve constitutive model calibrations and predictions.
@@
The remainder of this paper is organized as follows: First, the components of the ICC framework as applied to load path selection of a cruciform specimen in a biaxial load frame are presented (Sec. 2). Next, a discussion of the experimental setup, constitutive model, statistical methods and dimension reduction of full-field data, as well as all other details of the framework are described in detail (Sec. 3). The ICC framework is demonstrated on an exemplar problem to calibrate an elasto-plastic model using synthetically generated full-field DIC and global load data (Sec. 4). Finally, various ICC algorithmic decisions and considerations for future work are discussed (Sec 5). In summary, this work brings together several recent advancements in model calibration into one enhanced calibration framework to improve both the efficiency and results of constitutive model calibration.
2 ICC Framework
@@
The steps of the framework are illustrated in Fig. 1. Starting with plot (a), the experiment begins with an initial load step—deforming the cruciform specimen along one of the two axes while holding the other axis fixed (displacement = 0). DIC field data and global load data are collected in (b). The field data is then dimensionally reduced in (c) via principal component analysis (PCA) in order to make the algorithm more numerically stable. Furthermore, working in a reduced dimensional space aided in the construction of surrogate models which were pre-built. In (d), model calibration is performed via Bayesian inference, which yields a distribution that describes the parameter uncertainty. The expected information gain (EIG) is then optimized over the load step options (applying an increment of displacement along either of the two axes) to determine the next load step in (e). Finally, in order to efficiently perform the calibration with an expensive FE model, in (f), a surrogate replacement of the model is used (built off-line prior to entering the feedback loop) which is computationally cheap to evaluate. Additionally, in order to control the experiment, the infinite load path space is reduced to a finite load path tree—as seen in (f)—, which enables load step selection—details following. The feedback loop continues for some pre-determined number of cycles (equivalently, load steps) and yields a calibrated model with quantified uncertainty given the load path that is chosen. Each of the components which comprise the ICC framework in this setting are discussed in detail in Sec. 3.
@@
Figure 1: The ICC framework is shown for a cruciform specimen being deformed in an actively controlled biaxial load frame. An initial displacement is imposed on the cruciform along a single axis while keeping the other axis fixed (a). Full-field DIC and global load data are collected (b). The high dimensionality of the DIC data is reduced via PCA (c). Given the collected data and any prior knowledge on the model parameters, Bayesian inference is used to update the knowledge about the parameters with quantified uncertainty (d). The next load step is chosen as the one which has the greater EIG (e). Prior to entering the feedback loop, a surrogate model is constructed to approximate the expensive FE simulation at the locations where data may be collected and organized into a load path tree (f).
@@
In reality, the load path space from which the optimal load path is being selected is infinite when considering strain rate, step size, direction, etc. In order to enable the selection of an optimal load path in real-time, the decision space must be moved from infinite to finite dimensions. This is done by reducing the infinite-dimensional load path space of material deformation to a discrete space through the introduction of a load path tree.
@@
The same graph structure that was introduced in [27] is adopted here. The graph structure is depicted in Fig. 2, which is a binary tree that stands in for the biaxial loading of the cruciform specimen. The BOED workflow begins at the root node with the cruciform in its undeformed state. Each parent node has two children nodes that correspond to applying a fixed displacement increment along either Axis A (the horizontal axis) or Axis B (the vertical axis) of the cruciform while holding the alternate axis fixed. The ICC framework is not restricted to this tree structure, but it was chosen for this demonstration as a simple starting point. Other more complex, and perhaps more appropriate, options are possible and are discussed Sec.5.
@@
Figure 2: The load path tree reduces the infinite-dimensional load path space to finite dimensions. Here, a binary tree is shown which stands in for the deformation of a cruciform specimen in a biaxial load frame. The top node represents the specimen in its undeformed state. Each parent node has two children corresponding to applying an increment of displacement along either Axis A (horizontal) or Axis B (vertical) while holding the alternate axis fixed. In the example tree, four possible load paths (1-4) are shown, each consisting of two load steps.
@@
In summary, the ICC framework utilizes a combination of complex specimen geometry, load path optimization, and full-field DIC calibration data, leading to fewer experiments and increased information content for model calibration compared to traditional approaches. As a result, well-calibrated constitutive models with quantified parameter uncertainty (a natural output of BOED) supports decision-making at earlier stages in the design cycle.
@@ 3 Methods
In this section, the various methods and techniques utilized to perform material characterization and model calibration in the ICC framework are discussed. The model of the physical system is discussed in Sec. 3.1, with details of the model form in Sec. 3.1.1, the cruciform and finite element analysis in Sec. 3.1.2 and the synthetic data generation in Sec. 3.1.3. A description of dimension reduction of the DIC data through PCA is found in Sec. 3.2, and details of the surrogate model construction are found in Sec. 3.3. Bayesian techniques for inference and experimental design are discussed in Sec. 3.4.1 and Sec. 3.4.2, respectively.
@@ 3.1.1 Material Model Form
The final consideration in the selection of the model is the yield function. In the current study, multiaxial loadings are being leveraged enabling the consideration of non-quadratic and/or anisotropic yield functions. Multiple yield functions in those classes have been used to probe the response of various aluminum alloys and corresponding forms [28, 29, 30, 31, 32]. Each model has a different degree of complexity and corresponding number of calibration parameters. As this effort represents the first attempt at the novel ICC calibration approach, the relatively simple Hosford yield function [33] is considered to enable testing of the proposed scheme. While isotropic, the Hosford surface needs only a single calibration coefficient (the exponent, $a$) with different strengths in shear and tension. The single parameter enables keeping the dimensionality of the calibration optimization problem appropriate while introducing a term sensitive to the yield surface shape. Further, the Hosford model has been shown to reasonably reproduce various responses in aluminum alloys (e.g.  [28, 30]).
@@
in which $\sigma_{ij}$, $\mathbb{C}_{ijkl}$, $\varepsilon_{ij}$, and $\varepsilon_{ij}^{\text{p}}$ are the Cauchy stress, elastic stiffness (assumed isotropic), total strain, and plastic strain, respectively. The yield function, $g$, is then written,
@@
with $\kappa$ being the isotropic hardening variable (found via the consistency condition), $\phi$ the effective stress describing the shape of the yield function, and $\bar{\sigma}$ the flow stress capturing the size of the yield surface. The effective stress is taken to be that of Hosford and given as,
@@
In Eqn. (4), $\sigma_{y}$, $A$, and $n$ are the initial yield stress, exponential modulus, and exponential coefficient to be found via calibration. An associative flow rule is also utilized such that,
@@ 3.1.2 Cruciform Specimen
The key design parameters for the cruciform geometry used in the ICC framework were (1) a multi-axial/ complex state of stress, (2) multiple load-steps before failure, and (3) no yielding in the arms (i.e. failure would occur in the central gauge area instead of the arms). Many cruciform geometries exist in the literature [39, 40, 41, 42, 43, 44, 45, 46, 47, 48], but these geometries are not always in strict alignment with the above criteria. The cruciform for this work was designed according to the criteria presented in Marmos et al. [47], which capitalizes on using re-entrant radii and a thinned gauge area. The used geometry is shown notionally in Fig. 3, with complete dimensions included in Fig. 1 in A. Further details and considerations of the specimen design can be found in [49].
With an assumed eight-fold symmetry, one-eighth of the cruciform geometry was modeled, and symmetry displacement boundary conditions were applied to preserve computational resources. These boundary conditions consisted of a constant, uniform displacement at the distal nodes of each arm along the coordinate direction for Axes A or B while the other degrees of freedom were traction-free. The mesh was comprised of hexahedral elements with linear shape functions and a selective-deviatoric integration rule. The elements in the center region were considerably smaller than those in the arms.
The deformation of the specimen was driven by displacement boundary conditions applied at a constant rate, and the displacement increment was 0.25 mm per load step applied to each arm. For example, load path 1 in Fig. 2 consists of two load steps where Axis A (the horizontal arms of the cruciform) is displaced by 0.5 mm (0.25 mm twice) while Axis B (the vertical arms of the cruciform) is held fixed in displacement control. Given the symmetry boundary conditions, the displacement of each load step equates to 0.5 mm total extension along a given axis.
@@
Figure 3: Cruciform geometry and mesh. Left: Eight-fold symmetry was assumed such that only the opaque region was modeled. Displacement boundary conditions for Axis A (blue) and Axis B (red) were prescribed at the nodes at the end of each arm. The displacement surrogates predict values at nodes in the circular region (green). Right: The mesh was comprised of hexahedral elements of varying sizes.
An implicit, quasistatic finite-element analysis (FEA) was performed using Sierra/SM [35] with the given geometry and material model form (Sec. 3.1.1). An adaptive time-stepping algorithm was employed for the backward Euler time integrator used with the global equilibrium solution.The quantities of interest (QoIs) extracted from the FEA, which can in principal be computed from experimental measurements, include nodal displacements and resultant forces along each axis of the cruciform. The QoIs are returned for both the horizontal and vertical directional components and are denoted by X and Y, respectively. Thus the input-output relationship of the physical model is written as $g(\boldsymbol{\theta})=[\text{disp}_{i},\text{load}_{i}],i\in[X,Y]$. [^1] The horizontal and vertical directional components of displacement have traditionally been denoted by U and V, respectively; however, in order to avoid confusion with other notation in this paper, X and Y are used instead.

### v3

@@ Abstract
Accurate material characterization and model calibration are essential for computationally-supported high-consequence engineering decisions. Historically, characterization and calibration methods have (1) use simplified test specimen geometries and global data, (2) cannot guarantee that sufficient characterization data is collected for a specific model of interest, (3) use deterministic methods that provide best-fit parameter values with no uncertainty quantification, and (4) are sequential, inflexible, and time-consuming.
This work brings together several recent advancements into an improved workflow called Interlaced Characterization and Calibration (ICC) that advances the state-of-the-art in constitutive model calibration. The ICC paradigm (1) employs tools to efficiently use full-field data to calibrate high-fidelity material models, (2) aligns the data needed with the data collected by adopting an optimal experimental design protocol, (3) quantifies parameter uncertainty through Bayesian inference, and (4) incorporates these advancements into a quasi real-time feedback loop. The ICC framework is demonstrated here on the calibration of a material model using simulated full-field data for an aluminum cruciform specimen being deformed bi-axially. The cruciform is actively driven through the myopically preferred load path using Bayesian optimal experimental design, which selects load steps that yield the maximum expected information gain (EIG). Principal component analysis (PCA) is performed on the model predictions of full-field displacements, and fast surrogate models are built to approximate the input-output relationships of the expensive finite element model. The tools developed and demonstrated here show that high-fidelity constitutive models can be efficiently and reliably calibrated with quantified uncertainty, thus supporting credible decision-making and potentially increasing the agility of solid mechanics modeling by enabling utilization of computational simulations at earlier stages of the design cycle.SAND2024-15320O
@@
Bayesian optimal experimental design , load path selection , uncertainty quantification , constitutive model calibration , digital image correlation , principal component analysis , Gaussian process surrogate modeling
@@ 1 Introduction
Computational simulation is relied upon to understand complex engineering problems and provides critical information for decision-making. In finite element analysis (FEA) of solid mechanics, constitutive relationships describe a material’s response to external stimuli. These constitutive equations typically contain material-dependent parameters that are determined through material characterization and model calibration. The calibration and subsequent validation of these equations and parameters are crucial for accurate simulations and structural analyses, necessitating an emphasis on efficiency and accuracy in the calibration process to establish confidence in the engineering predictions.
Historically, material characterization and model calibration have frequently been performed sequentially and independently of each other, often requiring multiple experimental campaigns to gather adequate data for calibration [1]. This process typically involves assuming phenomenological features based on subject expertise, designing and executing characterization experiments based on those assumptions and calibrating a single material model deterministically. The accuracy of the calibration is not evaluated until the end of the process, and if the results are inadequate, the entire workflow may need to be repeated.
@@
In addition to the procedure, the type of data collected has contributed to the historical challenge of material characterization and model calibration. Traditionally, tests which target specific stress states (i.e. uniaxial tension, notched tension, top hat shear, etc.) have been used for calibration. These tests provide limited information per specimen, necessitating large test matrices which include various tests, sample geometries and cut directions in order to reveal complex material behaviors [2]. Further, such simplified states are often not reflective of real-world loading conditions. As a result of the described characterization and calibration workflow in tandem with traditional testing procedures, the model calibration process is elongated and expensive, thus delaying the support of computational simulation for design and engineering decisions.
@@
Recent advancements have contributed to significant improvements in constitutive model calibration to attempt to address these shortcomings. The use of digital image correlation (DIC) to collect rich full-field data through camera-based measurements has led to a paradigm shift in material characterization. Also known by the community as Material Testing 2.0 (MT2.0) [3], full-field DIC data collected from complex testing configurations (specimen shape plus loading) has greatly improved the information content that can be garnered from a single experiment for model calibration [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Moreover, researchers have developed several different methods—including Chebychev and Zernicke polynomials [15, 16, 17, 18, 19, 20] and various types of wavelets [21, 22, 23]—for decomposing the field data into a spectral domain in order to reduce the dimensionality of the data while retaining its rich information content.
The availability of full-field measurements has led to various methods for identification of mechanical properties and parameters from this data [24]. Updating techniques, such as Finite Element Model Updating (FEMU) and Constitutive Equation Gap Method (CEGM), compute displacement and stress fields using finite element analysis and optimize a cost function to align computed results with experimental data. In contrast, non-updating methods like the Virtual Fields Method (VFM) and Equilibrium Gap Method (EGM) derive stress fields from measured displacements based on a postulated constitutive model, using equilibrium equations to solve for the desired parameters. Finally data-driven approaches, such as demonstrated in [25], identify the strain-stress relationship of non-linear elastic materials without relying on a specific constitutive equation. These methods construct a database of strain-stress pairs from non-homogeneous strain fields, such as those obtained through DIC, to capture the material’s mechanical response.
Even with the improved characterization and parameter identification methods, selection of the proper experimental configuration, or design, which greatly influences the suitability of the data for model calibration, remains a challenging task. Thus, another area of improvement for model calibration is the development of procedures for optimal experimental design within the context of solid mechanics. In general, there are two different approaches which have been adopted: 1) optimization of the geometry of the specimen [26, 27, 28, 11], or 2) optimization of the load path of deformation [29, 30, 31]. Both approaches use various techniques to construct an experiment which will yield the most useful data for model calibration and, consequently, improve constitutive model calibrations and predictions.
@@
The remainder of this paper is organized as follows: First, the components of the ICC framework as applied to load path selection of a cruciform specimen in a biaxial load frame are presented (Sec. 2). Next, a discussion of the experimental setup, constitutive model, statistical methods and dimension reduction of displacement field data, as well as all other details of the framework are described in detail (Sec. 3). The ICC framework is demonstrated on an exemplar problem to calibrate an elasto-plastic model using synthetically generated full-field DIC and global load data (Sec. 4). Finally, various ICC algorithmic decisions and considerations for future work are discussed (Sec. 5). In summary, this work brings together several recent advancements in model calibration into one enhanced calibration framework to improve both the efficiency and results of constitutive model calibration.
2 Interlaced Characterization and Calibration (ICC) Framework
The ICC framework integrates material characterization and model calibration to enhance efficiency, reduce cost and accelerate the application of computational simulations in design. By prioritizing the collection of the most relevant data, ICC reduces uncertainties in material model parameters and increases confidence in model predictions to support informed decision-making.
@@ 2 Interlaced Characterization and Calibration (ICC) Framework
The steps of the framework are illustrated in Fig. 1. The solid arrows point to each successive step in the ICC framework, where data collection and model calibration occur in an iterative, quasi-real-time feedback loop. The dashed arrows stem from tasks that occur offline prior to entering the feedback loop and point to the steps where they are utilized. Starting with plot (a), the experiment begins with an initial load step—deforming the cruciform specimen along one of the two axes, which correspond to load step (LS) A or B, while holding the other axis fixed (displacement = 0). DIC field data and global load data are collected in (b). In (c), model calibration is performed via Bayesian inference, which yields a distribution that describes the parameter uncertainty. The expected information gain (EIG) is then calculated for the load step options (applying an increment of displacement along either of the two axes), and the load step that results in the greater EIG is determined to be the next load step in (d). The process then returns to (a), where the selected load step is taken. The feedback loop continues for some pre-determined number of cycles (equivalently, load steps) and yields a calibrated model with quantified uncertainty given the load path that is chosen.
Several tasks occur offline prior to entering the feedback loop. In order to control the experiment, the infinite load path space is reduced to a finite load path tree, as seen in (e), which enables load step selection in (d) and aids in the construction of surrogate models in (g). The FE model is evaluated at each node in the load path tree for various parameter combinations, and PCA is performed on the displacement fields (f). Finally, in order to efficiently perform steps (c) and (d) for the expensive finite element (FE) model, a surrogate replacement of the model is used that approximates the input-output relationship of the expensive model at each node in the load path tree and is computationally cheap to evaluate: The surrogate model output consists of global load values and PCA singular values for the dimensionally-reduced displacement fields. The displacement fields are reconstructed from the surrogate predictions of the PCA singular values in (h) before being used in the feedback loop. Each of the components which comprise the ICC framework in this setting is discussed in detail in Sec. 3.
@@
Figure 1: The ICC framework is shown for a cruciform specimen being deformed in an actively controlled biaxial load frame. Solid arrows point in the direction the ICC steps proceed, and dashed arrows stem from steps that happen offline prior to entering the ICC feedback loop and point towards steps where they are utilized. An initial displacement is imposed on the cruciform along a single axis, either the X (horizontal) or Y(vertical) axis corresponding to load step (LS) A or B, respectively, while keeping the other axis fixed (a). Full-field DIC and global load data are collected (b). Given the collected data and any prior knowledge on the model parameters, Bayesian inference is used to update the knowledge about the parameters with quantified uncertainty (c). The next load step is chosen as the one which has the greater EIG (d). Prior to entering the feedback loop, the infinite load path space is reduced to a finite representation through a load path tree (e). The FE model is evaluated at each node for various parameter combinations, and PCA is performed on the displacement fields (f). Surrogate models are constructed in the spectral domain to approximate the expensive FE simulation at each node in the load path tree (g), which are the locations where data may be collected. The displacement fields are reconstructed from the surrogate predictions of the PCA singular values before being used in the feedback loop (h), along with the predicted load.
@@
In reality, the load path space from which the myopically preferred load path is being selected is infinite when considering strain rate, step size, direction, etc. In order to enable the selection of the load path in real-time, the decision space must be moved from infinite to finite dimensions. This is done by reducing the infinite-dimensional load path space of material deformation to a discrete space through the introduction of a load path tree.
@@
The same graph structure that was introduced in [30] is adopted here. The graph structure is depicted in Fig. 2, which is a binary tree that stands in for the biaxial loading of the cruciform specimen. The BOED workflow begins at the root node with the cruciform in its undeformed state. Each parent node has two children nodes that correspond to applying a fixed displacement increment along either the horizontal axis or the vertical axis of the cruciform while holding the alternate axis fixed. The ICC framework is not restricted to this tree structure, but it was chosen for this demonstration as a simple starting point. Other more complex, perhaps more appropriate, options are possible and are discussed in Sec. 5. In this work, the two load step options are generically referred to as A and B, which correspond to applying an increment of displacement along the X (horizontal) or Y (vertical) axis, respectively.
@@
Figure 2: The load path tree reduces the infinite-dimensional load path space to finite dimensions. Here, a binary tree is shown which stands in for the deformation of a cruciform specimen in a biaxial load frame. The top node represents the specimen in its undeformed state. Each parent node has two children corresponding to applying an increment of displacement along either the X (horizontal) axis—load step (LS) A—or the Y (vertical) axis—LS B—while holding the alternate axis fixed. In the example tree, four possible load paths (1-4) are shown, each consisting of two load steps.
@@
In summary, the ICC framework utilizes a combination of complex specimen geometry, discovery of the preferred load path via BOED and full-field DIC calibration data, leading to fewer experiments and increased information content for model calibration compared to traditional approaches. As a result, well-calibrated constitutive models with quantified parameter uncertainty, a natural output of BOED, supports decision-making at earlier stages in the design cycle.
@@ 3 Methods
In this section, the various methods and techniques utilized to perform material characterization and model calibration in the ICC framework are discussed. The model of the physical system is discussed in Sec. 3.1, with details of the model form in Sec. 3.1.1, the cruciform and finite element analysis in Sec. 3.1.2 and the synthetic data generation in Sec. 3.1.3. A description of dimension reduction of the displacement field data through PCA is found in Sec. 3.2, and details of the surrogate model construction are found in Sec. 3.3. Bayesian techniques for inference and experimental design are discussed in Sec. 3.4.1 and Sec. 3.4.2, respectively.
@@ 3.1.1 Material Model Form
The final consideration in the selection of the model is the yield function. In the current study, multiaxial loadings are being leveraged enabling the consideration of non-quadratic and/or anisotropic yield functions. Multiple yield functions in those classes have been used to probe the response of various aluminum alloys and corresponding forms [32, 33, 34, 2, 35]. Each model has a different degree of complexity and corresponding number of calibration parameters. As this effort represents the first attempt at the novel ICC calibration approach, the relatively simple Hosford yield function [36] is considered to enable testing of the proposed scheme. While isotropic, the non-quadratic Hosford surface needs only a single calibration coefficient (the exponent, $a$) to describe different strengths in shear versus tension. The single parameter enables keeping the dimensionality of the calibration optimization problem appropriate while introducing a term sensitive to the yield surface shape. Further, the Hosford model has been shown to reasonably reproduce various responses in aluminum alloys (e.g.  [32, 34]).
@@
in which $\sigma_{ij}$, $\mathbb{C}_{ijkl}$, $\varepsilon_{ij}$ and $\varepsilon_{ij}^{\text{p}}$ are the Cauchy stress, elastic stiffness (assumed isotropic), total strain and plastic strain, respectively. The yield function, $g$, is then written,
@@
with $\kappa$ being the isotropic hardening variable (found via the consistency condition), $\phi$ the effective stress describing the shape of the yield function and $\bar{\sigma}$ the flow stress capturing the size of the yield surface. The effective stress is taken to be that of Hosford and given as,
@@
In Eqn. (4), $\sigma_{y}$, $A$ and $n$ are the initial yield stress, hardening modulus and hardening coefficient to be found via calibration. An associative flow rule is also utilized such that,
@@ 3.1.2 Cruciform Specimen
The key design parameters for the cruciform geometry used in the ICC framework were (1) a multi-axial/ complex state of stress, (2) multiple load-steps before failure and (3) no yielding in the arms (i.e. failure would occur in the central gauge area instead of the arms). Many cruciform geometries exist in the literature [42, 43, 44, 45, 46, 47, 48, 49, 50, 51], but these geometries are not always in strict alignment with the above criteria. The cruciform for this work was designed according to the criteria presented in Marmos et al. [50], which capitalizes on using re-entrant radii and a thinned gauge area. The geometry is shown notionally in Fig. 3, with complete dimensions included in Fig. S1 in the Supplementary Materials. Further details and considerations of the specimen design can be found in [52].
Exploiting symmetry about the three principal planes enables consideration of an eighth symmetry model as depicted in Fig. 3 to preserve computational resources. Corresponding symmetry displacement boundary conditions were used on these planes to enforce this condition. To achieve the desired experimental loadings, a constant and uniform displacement was applied at the distal nodes of each arm along the coordinate direction of either the X (horizontal) or Y (vertical) axis—corresponding to taking load step A or load step B in Fig. 2. Surfaces without displacement-specified boundary conditions were taken to be traction-free. The mesh was comprised of linear hexahedral elements with a selective-deviatoric integration rule. The elements in the center region were considerably smaller than those in the arms—Fig. 3.
The deformation of the specimen was driven by displacement boundary conditions applied at a constant rate, and the displacement increment was 0.25 mm per load step applied to each arm. For example, load path 1 in Fig. 2 consists of two sequential A load steps—denoted as load path AA—in which the end of each horizontal arm of the cruciform is displaced by 0.25 mm outward at each load step while the ends of the vertical arms are held in fixed in displacement control. Given the symmetry boundary conditions, the displacement of each load step equates to 0.5 mm total extension along a given axis.
@@
Figure 3: Cruciform geometry and mesh. The full geometry of the cruciform specimen is shown in (a) with the opaque region denoting the eighth-symmetry domain considered for computation. Displacement boundary conditions for the X axis and Y axis were prescribed at the nodes at the end of each arm. The displacement surrogates predict values at nodes in the circular region (green). The mesh of the computational domain used for the simulation is shown in (b) depicting the heterogeneous element size distribution.
An implicit, quasistatic FEA was performed using Sierra/SM [38] with the given geometry and material model form (Sec. 3.1.1). An adaptive time-stepping algorithm was employed for the backward Euler time integrator used with the global equilibrium solution. The quantities of interest (QoIs) extracted from the FEA, which can in principle be computed from experimental measurements, include nodal displacements and resultant forces along each axis of the cruciform. The QoIs are returned for both the X and Y directional components. Thus the input-output relationship of the physical model is written as $g(\boldsymbol{\theta})=[\text{disp}_{i},\text{load}_{i}],\,i\in[X,Y]$.